OpenStudy (anonymous):

5) WILL GIVE MEDAL AND FAN!! Write the product as a trinomial. (k – 1)(k – 4) a. k2 – 5k + 4 b. k2 – 5k – 4 c. k2– 4 d. k2 + 5k + 4

OpenStudy (alexandervonhumboldt2):

ok so use FOIL method

OpenStudy (alexandervonhumboldt2):

(k – 1)(k – 4)=k^2-4k-k+4=simplify

OpenStudy (anonymous):

Just use FOIL (First, Outer, Inner, Last): (k - 2)(k - 1) First: k * k = k² Outer: k * -1 = -k Inner: -2 * k = -2k Last: -2 * -1 = 2 Combine the terms: k² - k - 2k + 2 Answer: k² - 3k + 2

OpenStudy (anonymous):

@brae

OpenStudy (anonymous):

YAY :)

OpenStudy (anonymous):

wait so what answer would this be? sorry guys! @Hannah_Waller

OpenStudy (anonymous):

idk what i got is not in your choices :( im pretty sure i did it right though...

OpenStudy (anonymous):

ohhh... ummm @satellite73 @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

ok so i got A cause -4k-k=-5k and -1*-4=+4

OpenStudy (alexandervonhumboldt2):

i got A and it is correct

OpenStudy (anonymous):

thanks guys!! @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

np @Hannah_Waller just realized you were wrong

OpenStudy (anonymous):

haha its Ok dang so close haha :)